# A Basic Evaluation of Neural Networks Trained with the Error Diffusion Learning Algorithm

### v2

@@ 4.4 Analysis of Internal Representations
Figure 9 shows a heatmap of the synaptic weights connecting the input layer to the hidden layer within the EDLA network, specifically trained to classify the digit “zero” from handwritten digits. The heatmap reveals that the network learned meaningful features from the input data. Notably, the neurons in the hidden layer exhibit a strong response to specific patterns within the images. For instance, the weights connecting the positive neurons in the input layer to the positive neurons in the hidden layer (denoted as "P-P" connections) show high activation, shaping the digit “zero.” This result suggests that the EDLA method effectively enables neurons to capture and represent the relevant digit-specific features. Additionally, similar "zero-shaped" patterns also emerge in other synaptic connections.
Synaptic weight distributions in neural networks provide valuable insights into the learning behavior of the network and features extracted from the input data. To explore this further, we analyze the distribution of weights across the EDLA network. Figure 10 shows a histogram representing the synaptic weights within the trained EDLA network. The distribution appears to be predominantly centered around small magnitudes, indicating that most weights are relatively small. However, a distinct spread is observed with a smaller subset of synapses exhibiting large positive or negative values. This broad distribution suggests that the EDLA update unilaterally increase the absolute values of the weights while balancing the excitatory and inhibitory weights.
Figure 9: A heatmap of the synaptic weights connecting the input layer to the first hidden layer in the EDLA network is trained to classify the digit “zero” in the Digits dataset. Each subplot shows the weight patterns for individual connection patterns, labeled according to connection types: P-P (positive input to positive hidden sublayers), N-P (negative input to positive hidden sublayers), P-N (positive input to negative hidden sublayers), and N-N (negative input to negative hidden sublayers). Color gradients denote weight values, with yellow indicating strongly positive weights and purple indicating strongly negative weights.
@@
5 Conclusion
@@
The Error Diffusion Learning Algorithm (EDLA) provides a biologically inspired alternative to conventional backpropagation methods for training neural networks and is characterized by the diffusion of a global error signal across network layers. This mechanism is analogous to reward-based learning processes observed in biological neural systems. Moreover, the EDLA mirrors neuromodulatory systems, such as the dopaminergic system, which projects broadly across cortical and subcortical areas, disseminating their signals globally (Roelfsema and Holtmaat, 2018). In contrast to standard MLPs, the EDLA incorporates structured interactions between neurons with excitatory and inhibitory synapses. This reflects the biologically observed balance between these types of synapses.
@@
However, EDLA is inherently restricted by its single-output architecture, which limits its applicability in multiclass classification tasks. A practical solution involves parallelizing multiple EDLA networks, although this approach increases computational costs. This parallelized architecture mirrors the modular structure of biological neural systems in which specialized circuits process distinct features or tasks independently. In such a configuration, each EDLA network separately optimizes its own output according to a unique global error signal to that network. This modular structure aligns with distributed processing mechanisms observed in biological neural networks.
@@
The EDLA demonstrates efficacy across various tasks, including parity checks, regression, and image classification. Experimental evaluations indicate that the EDLA achieves a satisfactory performance across various datasets. Notably, the ReLU AF achieves a performance comparable to that of the sigmoid function originally proposed by Kaneko (Kaneko, 1999), where the ReLU often converges more quickly. However, because the ReLU lacks an upper bound, it can produce huge and potentially unstable weight updates under the Hebbian-like learning dynamics employed by the EDLA. Moreover, while increasing the number of neurons in a hidden layer improves the performance and accelerates convergence, excessive network depth or an overly large number of hidden neurons may degrade the performance because of unstable weight updates. A reduction in the learning rate can mitigate these instabilities.
In contrast to backpropagation, which requires error signals to be propagated backward through each intermediate layer, EDLA relies solely on a single global error signal combined with local neural activities. This approach presents potential advantages for developing more biologically plausible neural network training techniques. Thus, the EDLA will represent a significant step toward bridging the gap between biological realism and artificial neural network methodologies.

### v3

@@ 4.4 Analysis of Internal Representations
Figure 9 shows a heatmap of the synaptic weights connecting the input layer to the hidden layer within the EDLA network, specifically trained to classify the digit “zero” from handwritten digits. The heatmap reveals that the network learned meaningful features from the input data. Notably, the neurons in the hidden layer exhibit a strong response to specific patterns within the images. For instance, the weights connecting the positive neurons in the input layer to the positive neurons in the hidden layer (denoted as "P-P" connections) show high activation, shaping the digit “zero.” This result suggests that the EDLA method effectively enables neurons to capture and represent the relevant digit-specific features. Additionally, similar “zero-shaped” patterns also emerge in other synaptic connections.
Synaptic weight distributions in neural networks provide valuable insights into the learning behavior of the network and features extracted from the input data. To explore this further, we analyze the distribution of weights across the EDLA network. Figure 10 shows a histogram representing the synaptic weights within the trained EDLA network. The distribution appears to be predominantly centered around small magnitudes, indicating that most weights are relatively small. However, a distinct spread is observed with a smaller subset of synapses exhibiting large positive or negative values. This broad distribution suggests that the EDLA update unilaterally increases the absolute values of the weights while balancing the excitatory and inhibitory weights.
Figure 9: A heatmap of the synaptic weights connecting the input layer to the first hidden layer in the EDLA network, trained to classify the digit “zero” in the Digits dataset. Each subplot shows the weight patterns for individual connection patterns, labeled according to connection types: P-P (positive input to positive hidden sublayers), N-P (negative input to positive hidden sublayers), P-N (positive input to negative hidden sublayers), and N-N (negative input to negative hidden sublayers). Color gradients denote weight values, with yellow indicating strongly positive weights and purple indicating strongly negative weights.
@@
4.5 Stabilization of Learning Dynamics in EDLA with ReLU Activation Function
Deeper EDLA networks employing the ReLU AF often exhibit severe performance degradation when the number of hidden layers (depth) increases. This degradation arises from the unbounded nature of the ReLU AF, which can generate excessively large activations and thereby produce excessively large and unstable weight updates during training. To mitigate these effects, two stabilization techniques are introduced: the use of small-scale initial weights and RMS normalization of neuronal activations (Sec. 2.7). The former limits the magnitude of initial activations, while the latter normalizes neuronal pre-activations based on their root-mean-square (RMS) values. These mechanisms jointly stabilize the learning dynamics by suppressing excessive activations and preventing abrupt weight updates.
Figure 11 presents the results of EDLA networks with the ReLU AF under these stabilization strategies across a range of network widths and depths for both regression and image classification tasks. For regression (Airfoil and Concrete datasets) and classification (MNIST and CIFAR-10 datasets), the learning rates are set to 0.0001 and 0.1, respectively. The weights are drawn from uniform distributions over $0,\mathrm{IS}]$ and $\mathrm{IS},0]$ for synapses connecting same-type and different-type neurons, respectively. The initial scale (IS) was 0.0001 for the small-scale variant, and 1.0 for both the standard and RMS normalization. RMS normalization is applied to each layer except the input and output layers. The mini-batch sizes are 64 for regression and 128 for classification.
For regression tasks, both small initial weights and RMS normalization suppress extreme or NaN (not-a-number) MAE values that arise in deeper networks, leading to stable training. For classification tasks, small initial weights do not significantly influence the accuracy of shallow networks (1–2 layers), but in deeper configurations (4 layers) on MNIST, they tend to degrade accuracy slightly. In contrast, RMS normalization improves accuracy in deeper architectures, particularly for the CIFAR-10 dataset with larger hidden dimensions. These findings indicate that small initial weights and RMS normalization are effective in preventing instability in deeper EDLA networks with the ReLU AF, stabilizing training without compromising general performance.
Figure 11: Performance comparison of EDLA with ReLU AF using two stabilization techniques: small initial weights (IS = 0.0001, left column) and RMS normalization (right column). The plots show performance across varying layer configurations (1, 2, and 4 layers) and neuron counts. Top two rows (Regression): MAE on Airfoil and Concrete datasets (LR = 0.0001). Bottom two rows (Classification): Accuracy on MNIST and CIFAR-10 datasets (LR = 0.1). Results are averaged over four independent trials.
The observed improvement in deeper EDLA networks likely arises from the suppression of exploding activations by these stabilization techniques. The instability of ReLU-based EDLA networks can primarily be attributed to two factors: (i) exploding activations, and (ii) the so-called dying ReLU phenomenon, in which neurons become inactive and cease to contribute to learning. The former refers to the uncontrolled growth of activations during forward propagation, which induces large Hebbian-like weight updates under EDLA’s learning rule. This effect becomes increasingly pronounced as the depth of the network grows, resembling the well-known exploding gradient phenomenon in recurrent or very deep networks [31, 32]. The latter, known as the dying ReLU problem [33, 34], occurs when neurons become inactive (outputting zero for all inputs), preventing them from contributing to learning. In practice, these two effects interact: large activations in early training produce extreme weight updates, which then push many neurons into permanently inactive states, further deteriorating the network’s representational capacity.
To elucidate the effects of these stabilization methods and clarify the cause of instability, several internal metrics were monitored. Test MAE was evaluated on the test set to track generalization performance. Concurrently, several internal dynamics were monitored on the training data, including the network-wide maximum and mean activations, per-layer activation maxima, the network-wide maximum and mean absolute weight updates, and the global dead-unit ratio (the percentage of inactive neurons). In this study, a neuron was considered dead if its activation remained below the fixed threshold of 0.01 for more than 95% of training samples. Figure 12 compares these metrics across three EDLA variants: the baseline EDLA with the ReLU AF, the EDLA with small-scale initialization (IS = 0.0001), and the EDLA with RMS normalization. The initial scale (IS) was set to 1.0 for both the baseline EDLA and the EDLA with RMS normalization. This internal analysis was conducted on the Concrete regression task using a network with four hidden layers of 32 neurons each, a learning rate of 0.0001, and a mini-batch size of 64.
As shown in Figure 12A, the baseline EDLA exhibits large and unstable test MAE values throughout training, indicating poor convergence and sensitivity to parameter updates. In contrast, both small initial weights and RMS normalization suppress large fluctuations in MAE, leading to smoother and smaller test errors. The small-scale initialization achieves the lowest MAE overall, while RMS normalization provides similarly stable and improved performance. Figures 12B and 12C show that the baseline network experiences a rapid increase in the maximum activation magnitude early in training. The small-scale initialization moderates this activation growth, and RMS normalization stabilizes activations at moderate levels throughout training. Interestingly, the RMS-normalized network maintains higher mean activations than the other variants, even when its maximum activations are well controlled, suggesting that RMS normalization prevents neuron inactivity while maintaining balanced activation levels. This interpretation is supported by the lower dead-unit ratio observed in Figure 12I.
The per-layer activation trends in Figures 12D–F reveal that deeper layers tend to exhibit larger activations due to the cumulative effect of the unbounded ReLU activation, while both stabilization techniques attenuate this amplification across all layers. Figures 12G and 12H further show that the baseline EDLA undergoes excessively large weight updates in the early stages of training. In the standard EDLA, although the mean weight update is smaller than that of the stabilized variants, the weights of many neurons remain nearly unchanged, while a few experience extremely large updates, leading to unstable learning dynamics. Consequently, the global dead-unit fraction shown in Figure 12I reveals that the baseline model suffers from a high proportion of inactive neurons. Many neurons become permanently dead early in training, severely limiting the network’s representational capacity. In contrast, both stabilization methods maintain smaller and more consistent weight update magnitudes across epochs. Both stabilization methods substantially reduce the occurrence of dead units, with RMS normalization showing the most pronounced effect. These results indicate that the stabilization techniques effectively regulate activation and weight dynamics.
These results collectively demonstrate that the instability of ReLU-based EDLA networks arises from the combined effects of exploding activations and dead ReLU units. Small initial weights alleviate early activation explosions, whereas RMS normalization further stabilizes learning by controlling both activation magnitudes and update scales. Importantly, RMS normalization maintains sufficient neuron activity and prevents the collapse of representational diversity. Overall, these findings indicate that appropriate control of activation magnitudes is essential for stable training in EDLA networks. The effectiveness of RMS normalization in particular parallels the success of normalization techniques such as Batch Normalization [26] and Layer Normalization [25], highlighting that regulation of internal activation dynamics is a key factor for achieving stable and efficient learning in non-standard architectures such as EDLA.
Figure 12: Analysis of internal dynamics for three EDLA-ReLU variants on the Concrete regression task ($n_{\mathrm{hidden}}=32$, $L=4$, LR = 0.0001). Variants are: ReLU (scale = 1.0) (baseline EDLA), ReLU (scale = 0.0001) (EDLA with the small initial scale), and RMS Norm + ReLU (scale = 1.0) (EDLA with RMS normalization). Metrics shown: (A) Test MAE; (B) network-wide maximum activation; (C) network-wide mean activation; (D–F) per-layer post-activation maxima (indices 1–4 correspond to hidden layers); (G) network-wide maximum absolute weight update; (H) network-wide mean absolute weight update; (I) global dead-unit fraction. Metric (A) was evaluated on the test set, while metrics (B)–(I) were evaluated on the training set. (D–H) are plotted on a log10 scale ($\varepsilon=10^{-12}$ added to avoid $\log(0)$).
4.6 Comparison with multilayer perceptron trained with backpropagation
To benchmark the performance of the EDLA network, it is compared with a standard multilayer perceptron (MLP) trained using backpropagation. The MLP architecture mirrors that of the EDLA network, consisting of an input layer, hidden layers, and an output layer. The neurons in the hidden layers utilize either the sigmoid or ReLU AFs, while the output layer employs the identity AF for regression tasks and the sigmoid AF for classification tasks. The MLP is trained using the AdamW optimizer (without weight decay), with default parameters except for the learning rate. The learning rates are set to 0.001 for the MLP with the sigmoid AF and 0.0001 for the MLP with the ReLU AF in regression tasks. For classification tasks, the learning rates are set to 0.01 for the MLP with the sigmoid AF and 0.001 for the MLP with the ReLU AF. Training is conducted for 500 epochs with a mini-batch size of 64 for regression tasks and 128 for classification tasks. For CIFAR-10, training is conducted for 1000 epochs with a mini-batch size of 128. The training configurations of EDLA variants are the same as in previous sections.
Table 1 presents the performance comparison across four representative datasets. In the “Model (Act)” column, the activation function and stabilization variant are indicated in parentheses. “IS” refers to the initial scale parameter used for weight initialization, and “RMS” indicates the application of RMS normalization to the pre-activations. For example, “EDLA (ReLU, IS=0.0001)” represents the ReLU-based EDLA with a small initial scale of 0.0001, and “EDLA (ReLU+RMS)” represents the ReLU-based EDLA incorporating RMS normalization.
For a fair comparison, the parameter counts of EDLA and MLP in each configuration were matched or approximated as closely as possible. The results confirm that the backpropagation-trained MLP is a more powerful optimizer, consistently achieving superior performance. However, the magnitude of this performance gap is highly dependent on task complexity and network depth. On simpler, shallower tasks, EDLA is highly competitive. On the 1-layer MNIST task, for instance, the best EDLA variant (Sigmoid) achieves 0.973 accuracy, only slightly trailing the MLP’s 0.981. In contrast, the gap widens significantly on deeper, more complex tasks. On the 4-layer Airfoil regression task, the MLP (ReLU) achieves an MAE of 1.599, whereas the best-performing stabilized EDLA (ReLU, Small IS) only reaches 6.381. This highlights a key finding: while the stabilization techniques (Sec. 4.5) successfully prevented the catastrophic failure of deep EDLA-ReLU networks (which produced MAE $10^{30}$), they were not sufficient to close the fundamental optimization gap with a standard backpropagation-trained MLP.
This performance gap, however, is expected and illustrates the fundamental trade-offs of the EDLA framework. Backpropagation achieves its high performance by propagating exact, layer-specific error gradients, allowing for fine-grained optimization across the entire network. EDLA, by design, sacrifices this optimization power for biological plausibility, relying instead on an approximated, single global error signal diffused uniformly to all layers. Furthermore, this gap is compounded on multi-class tasks (MNIST, CIFAR-10) by EDLA’s multi-head architecture. An MLP uses a shared representational body and a single final layer for all classes. In contrast, the EDLA implementation must train $K$ independent networks, one for each output class. This approach not only requires significantly more parameters (as analyzed in Sec. 2.6) but also prevents the network from learning shared, general-purpose features across classes, likely degrading performance on high-dimensional, multi-class problems.
On the CIFAR-10 dataset, the MLP’s performance is also relatively low, but this result is not irregular for a simple MLP architecture. Simple MLPs with ReLU show about 50% accuracy on CIFAR-10. For example, a simple MLP with ReLU activation, six layers, and 1024 neurons per layer achieves 54.2% accuracy with Layer Normalization on CIFAR-10 [35], while a two-layer MLP (width 512, ReLU, and 0.2 dropout) achieves 54–58% accuracy [36]. Thus, the MLP’s performance in the experiments is consistent with the findings in the literature, confirming that both EDLA and traditional MLPs encounter similar architectural limitations when applied to high-dimensional image data without convolutional priors.
Table 1: Performance comparison (MNIST, CIFAR-10, Airfoil, Concrete). For classification tasks the metric is accuracy; for regression tasks the metric is mean absolute error (MAE). The best performance for each dataset and architecture class (e.g., 4-layer) is highlighted in bold. Note that the parameter counts for EDLA and MLP include biases.
Dataset Metric Model (Act) Arch (L x N) Params Value (Mean)
Airfoil MAE EDLA (Sigmoid) 1 x 256 7.2k 3.828
MAE EDLA (ReLU) 1 x 256 7.2k 3.439
MAE EDLA (ReLU, IS=0.0001) 1 x 256 7.2k 3.751
MAE EDLA (ReLU+RMS) 1 x 256 7.2k 3.259
MAE MLP (Sigmoid) 1 x 1024 7.2k 2.411
MAE MLP (ReLU) 1 x 1024 7.2k 2.739
MAE EDLA (Sigmoid) 4 x 256 1.58M 9.072
MAE EDLA (ReLU) 4 x 256 1.58M $1.90\times 10^{31}$
MAE EDLA (ReLU, IS=0.0001) 4 x 256 1.58M 6.381
MAE EDLA (ReLU+RMS) 4 x 256 1.58M 20.941
MAE MLP (Sigmoid) 4 x 768 1.78M 5.842
MAE MLP (ReLU) 4 x 768 1.78M 1.599
Concrete MAE EDLA (Sigmoid) 1 x 256 10.2k 7.832
MAE EDLA (ReLU) 1 x 256 10.2k 7.511
MAE EDLA (ReLU, IS=0.0001) 1 x 256 10.2k 8.357
MAE EDLA (ReLU+RMS) 1 x 256 10.2k 6.271
MAE MLP (Sigmoid) 1 x 1024 10.2k 4.349
MAE MLP (ReLU) 1 x 1024 10.2k 4.630
MAE EDLA (Sigmoid) 4 x 256 1.59M 7.986
MAE EDLA (ReLU) 4 x 256 1.59M $9.84\times 10^{31}$
MAE EDLA (ReLU, IS=0.0001) 4 x 256 1.59M 19.814
MAE EDLA (ReLU+RMS) 4 x 256 1.59M 14.236
MAE MLP (Sigmoid) 4 x 768 1.78M 13.722
MAE MLP (ReLU) 4 x 768 1.78M 3.084
MNIST Accuracy EDLA (Sigmoid) 1 x 16 503.1k 0.973
Accuracy EDLA (ReLU) 1 x 16 503.1k 0.969
Accuracy EDLA (ReLU, IS=0.0001) 1 x 16 503.1k 0.970
Accuracy EDLA (ReLU+RMS) 1 x 16 503.1k 0.971
Accuracy MLP (Sigmoid) 1 x 512 407.1k 0.981
Accuracy MLP (ReLU) 1 x 512 407.1k 0.981
Accuracy EDLA (Sigmoid) 4 x 64 3.00M 0.968
Accuracy EDLA (ReLU) 4 x 64 3.00M 0.949
Accuracy EDLA (ReLU, IS=0.0001) 4 x 64 3.00M 0.942
Accuracy EDLA (ReLU+RMS) 4 x 64 3.00M 0.962
Accuracy MLP (Sigmoid) 4 x 1024 3.96M 0.980
Accuracy MLP (ReLU) 4 x 1024 3.96M 0.980
CIFAR-10 Accuracy EDLA (Sigmoid) 1 x 32 3.93M 0.421
Accuracy EDLA (ReLU) 1 x 32 3.93M 0.410
Accuracy EDLA (ReLU, IS=0.0001) 1 x 32 3.93M 0.364
Accuracy EDLA (ReLU+RMS) 1 x 32 3.93M 0.387
Accuracy MLP (Sigmoid) 1 x 1024 3.16M 0.471
Accuracy MLP (ReLU) 1 x 1024 3.16M 0.536
Accuracy EDLA (Sigmoid) 4 x 16 2.03M 0.193
Accuracy EDLA (ReLU) 4 x 16 2.03M 0.216
Accuracy EDLA (ReLU, IS=0.0001) 4 x 16 2.03M 0.243
Accuracy EDLA (ReLU+RMS) 4 x 16 2.03M 0.353
Accuracy MLP (Sigmoid) 4 x 512 2.37M 0.451
Accuracy MLP (ReLU) 4 x 512 2.37M 0.541
5 Conclusion and Discussion
@@
The Error Diffusion Learning Algorithm (EDLA) provides a biologically inspired alternative to conventional backpropagation methods for training neural networks and is characterized by the diffusion of a global error signal across network layers. This mechanism is analogous to reward-based learning processes observed in biological neural systems. Moreover, EDLA mirrors neuromodulatory systems, such as the dopaminergic system, which projects broadly across cortical and subcortical areas, disseminating their signals globally [12]. In contrast to traditional MLPs, EDLA incorporates structured interactions between neurons with excitatory and inhibitory synapses. This reflects the biologically observed balance between these types of synapses.
@@
However, EDLA is inherently constrained by its single-output architecture, limiting its applicability to multiclass classification tasks. A practical solution involves parallelizing multiple EDLA networks, although this approach increases computational costs. This parallelized architecture mirrors the modular structure of biological neural systems in which specialized circuits process distinct features or tasks independently. In such a configuration, each EDLA network separately optimizes its own output according to a unique global error signal to that network. This modular structure aligns with distributed processing mechanisms observed in biological neural networks.
@@
EDLA demonstrates efficacy across various tasks, including parity-check, regression, and image classification. Experimental evaluations indicate that the EDLA achieves satisfactory performance across various datasets. Notably, the ReLU AF achieves a performance comparable to that of the sigmoid function originally proposed by Kaneko [9], where the ReLU often converges more quickly. The faster convergence of ReLU networks observed here replicates earlier findings showing that the ReLU AF accelerates learning in neural networks [37].
In contrast, the experiments in this study consistently showed that increasing EDLA’s network depth, unlike standard deep learning, does not reliably improve performance and can even lead to degradation (Figs. 6, 7, 8). We hypothesize this stems from EDLA’s unique update mechanism based on a single global error signal. Unlike backpropagation, where gradients can vanish in deep layers due to the chain product of activation function derivatives [38, 32], EDLA’s global signal ensures all weights receive an update signal. While this avoids the classical vanishing gradient problem caused by the chain product of derivatives, it does not guarantee effective credit assignment for deep hierarchies. Updates driven by the single scalar error $d$ might not be optimal for each specific layer’s needs. Furthermore, changes induced in lower-layer activations propagate upwards and can be amplified through successive layers. This amplification becomes particularly severe (“exploding activations”) when using unbounded activation functions like ReLU. However, even with bounded functions like sigmoid, the cumulative effect of layer-by-layer amplification driven by a non-specific global signal might interfere with learning the complex, hierarchical features typically expected from deeper networks, potentially explaining the lack of performance improvement or degradation. This amplification effect is inherently less pronounced in shallower networks. Consequently, for EDLA, increasing network width appears to be a more effective and stable strategy for improving performance than increasing depth.
A key finding was the depth-related instability observed, especially with the ReLU activation function. This instability manifested severely in deep ReLU regression networks, causing catastrophic failures (e.g., NaN MAE). This failure mode is likely exacerbated by the regression task’s linear output layer: exploding activations from the final hidden layer directly produce an extremely large output $y$, which in turn generates an enormous global error signal d=tyd=t-y. This huge error signal then catastrophically amplifies weight updates throughout the entire network. This depth-related instability primarily stems from the unbounded nature of the ReLU AF itself, a phenomenon observed in standard deep networks as well [31, 32]. As signals propagate through multiple layers, ReLU’s lack of saturation allows activation magnitudes to grow excessively (“exploding activations”). While this is a general challenge in deep ReLU networks, it becomes particularly problematic within the EDLA framework. EDLA’s Hebbian-like learning rule translates these large activations directly into excessively large weight updates. The single global error signal, while perhaps not the primary initiator of the explosion, likely contributes to propagating instability across the network once large activations emerge.
Crucially, the internal dynamics analysis (Figure 12) provides direct quantitative evidence for this cascade in EDLA, but reveals a mechanism for vanishing updates different from classical vanishing gradient. The baseline deep ReLU network exhibited exploding maximum activations (Fig. 12B) and extreme initial weight updates (Fig. 12G, H), immediately followed by a dramatic rise in the dead neuron fraction to over 80% (Fig. 12I). This strongly indicates that performance collapse stems from ReLU-driven exploding activations, exacerbated by EDLA’s update rule, triggering widespread neuron death, which is itself a primary source of vanishing weight updates, as dead units provide zero update signals. In contrast, the saturating sigmoid function inherently limits activation magnitudes and thus update sizes, explaining its greater stability in EDLA.
It was demonstrated that this instability can be effectively managed. While smaller learning rates and initial weights help, RMS normalization proved particularly robust, controlling extreme activations while crucially maintaining neuron activity (low dead unit fraction, Fig. 12I), thus enabling stable learning in deeper ReLU networks.
Despite successful stabilization, comparison with backpropagation-trained MLPs (Table 1) revealed a persistent performance gap, widening with network depth and task complexity. Although EDLA achieved highly competitive results in shallow configurations, MLPs consistently performed better overall. This underscores EDLA’s fundamental trade-off: sacrificing the optimization power of exact, layer-specific gradients for the biological plausibility of an approximated, global error signal. The multi-head architecture likely exacerbates this gap in multi-class settings by preventing shared feature learning.
Nonetheless, EDLA remains a significant contribution towards biologically realistic learning algorithms. Its reliance on local activity and a global signal holds promise for neuromorphic computing and insights into brain function.
Future work could explore several avenues to enhance EDLA’s capabilities. First, integrating eligibility traces could enable single EDLA networks to handle multiclass classification tasks, reducing the need for parallel architectures. Eligibility traces could allow neurons to maintain a memory of past activations, facilitating more complex decision boundaries. However, determining which neurons should maintain these traces remains an open question. Second, investigating hybrid architectures that combine EDLA with backpropagation in a complementary manner might leverage the strengths of both approaches. For instance, EDLA might be employed for initial coarse learning phases, followed by fine-tuning via backpropagation. Finally, applying EDLA to other architectures such as recurrent and convolutional neural networks [39, 40, 41] could extend its applicability to temporal and spatial data domains, respectively.